OpenStudy (anonymous):

find the area of a rectangle widow measuring 3/4 meter by 5/8 meter?

OpenStudy (bahrom7893):

A = (3/4)*(5/8) = 15/24 (m^2)

OpenStudy (bahrom7893):

woops sorry A = (3/4)*(5/8) = 15/32 (m^2)

OpenStudy (bahrom7893):

4*8 is 32 not 24.. Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)